### v1

# SPRec: Leveraging Self-Play to Debias Preference Alignment for Large Language Model-based Recommendations

Abstract.
Large language models (LLMs) have attracted significant attention in recommendation systems. Current LLM-based recommender systems primarily rely on supervised fine-tuning (SFT) to train the model for recommendation tasks. However, relying solely on positive samples limits the model’s ability to align with user satisfaction and expectations. To address this, researchers have introduced Direct Preference Optimization (DPO), which explicitly aligns recommendations with user preferences using offline preference ranking data. Despite its advantages, our theoretical analysis reveals that DPO inherently biases the model towards a few items, exacerbating the filter bubble issue and ultimately degrading user experience. In this paper, we propose SPRec, a novel self-play recommendation framework designed to mitigate over-recommendation and improve fairness without requiring additional data or manual intervention. In each self-play iteration, the model undergoes an SFT step followed by a DPO step, treating offline interaction data as positive samples and the predicted outputs from the previous iteration as negative samples. This effectively re-weights the DPO loss function using the model’s logits, adaptively suppressing biased items. Extensive experiments on multiple real-world datasets demonstrate SPRec’s effectiveness in enhancing recommendation accuracy and addressing fairness concerns.
Large Language Model-based Recommendation, Homogeneity Issue, Fairness, Self-play, Direct Preference Optimization
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
1. Introduction
Recently, large language models (LLMs) have demonstrated significant success across numerous domains, showcasing advanced capabilities in learning, reasoning, and generalizing to downstream tasks (Zhao et al., 2023; Wang et al., 2024a; Dai et al., 2024). In the field of recommender systems, there has been growing interest in leveraging the potential of LLMs in novel ways (Wu et al., 2024c). One prominent approach involves positioning LLMs as the central recommendation backbone, utilizing users’ past interactions and current needs to generate personalized item recommendations (Bao et al., 2023a; Wei et al., 2024). Compared to traditional methods, LLM-based recommendation systems (LRSs) offer distinct advantages, including a deeper contextual understanding and the flexibility to dynamically adapt to users’ evolving preferences.
Figure 1. Homogeneity issues in LLM-based recommendation results caused by token-level and item-level biases.
To enable LLMs to learn collaborative filtering signals and effectively perform specific recommendation tasks, a prevalent strategy is to fine-tune pre-trained LLMs via Supervised Fine-Tuning (SFT) on offline recommendation datasets (Bao et al., 2023a, b; Liao et al., 2024b; Zhang et al., 2023b). This approach allows LLMs to efficiently internalize user preferences within the offline data by adjusting their parameters to align with the recommendation task. Building on the foundation of SFT, recent research has adopted Direct Preference Optimization (DPO) (Rafailov et al., 2024) to further refine user preferences (Chen et al., 2024b; Bai et al., 2024; Liao et al., 2024a). While SFT relies solely on desirable answers, DPO incorporates both chosen and rejected response pairs, allowing the LLM to learn user ranking preferences and develop a more nuanced understanding of fine-grained, personalized information. This approach aligns with the common practice in recommendation models, which utilize both positive and negative samples to train the model effectively (Chen et al., 2023b; Shi et al., 2023).
Despite these advancements, we identify that employing DPO to align user preferences in recommender systems inherently introduces significant biases due to its underlying mechanisms. These biases can lead to serious homogeneity issues, where LLMs produce items with similar names or content. Fig. 1 provides a simplified visualization of how token-level and item-level biases manifest in the Top-$K$ movie recommendation results. Token-level biases occur due to the fact that LLMs generate item names in a tokenized fashion. Since LLMs are usually tuned to maximize the likelihood of target tokens, it can result in an overemphasis on items composed of more common tokens. For instance, movies with common words like “the” in their titles may be disproportionately recommended, regardless of their actual relevance to user preferences. At the item level, biases can emerge from multiple factors, particularly after fine-tuning, where LLMs may disproportionately recommend popular items, such as the Batman film series. This phenomenon can lead to the creation of filter bubbles (Gao et al., 2023a; Wang et al., 2022), where users are repeatedly exposed to a narrow range of highly popular content, ultimately detracting from the diversity of their recommendations and the overall user experience.
Some research has been conducted on bias and unfairness issues in LRSs (Dai et al., 2024; Zhang et al., 2023a; Wang et al., 2024b; Gallegos et al., 2024; Sakib and Das, 2024a). Dai et al. (2024) provided a comprehensive overview of the various types of biases that emerge across different stages of these models and outlined strategies to mitigate them. For example, Jiang et al. (2024) proposed re-weighting the fine-tuning loss for each item and re-ranking the generated results to ensure equal treatment of genre groups. Similarly, Bao et al. (2024) adjusted the LLM decoding process by removing the length normalization term and incorporating predictions from a text-free model, helping to reduce amplification bias and address homogeneity issues. However, these existing methods either require carefully crafted rules or rely heavily on external knowledge, limiting their applicability in general recommendation systems.
To this end, we propose a self-play recommendation tuning framework, SPRec, to adaptively suppress biases and improve fairness in LRSs without the need for additional information or expert knowledge. The core idea of SPRec is very straightforward: each tuning iteration begins with an SFT round using positive samples from offline data, followed by a DPO round still treating SFT data as positive samples and the predicted outputs from the previous iteration as negative samples. The philosophy is to let the model “play” with its own output by re-weighting the DPO loss function based on the predictions. Hence, the higher an item ranks in the model’s predictions, the more it is penalized. Meanwhile, the SFT process reinforces the ranking of positive items. Over time, this self-play learning process adaptively suppresses all undesirable items (biases) while maintaining alignment with positive samples. Extensive experiments on public datasets demonstrate that SPRec effectively enhances both accuracy and fairness metrics, showcasing its potential as a practical and efficient solution for LRSs.
The main contributions of this paper are as follows:
  • We provide an analysis that shows current LRSs tuned through DPO inevitably exhibit biases due to their underlying learning mechanisms, leading to the homogeneity issue.
  • We propose a self-play recommendation tuning framework, SPRec, which addresses these biases and improves fairness without the need for external knowledge.
  • Experiments validate SPRec’s effectiveness in improving accuracy, diversity, and fairness, with ablation studies attributing the improvements to the self-play negative samples.
2. Related work
We provide a brief overview of LLM-based recommender systems and their associated bias issues, followed by an introduction to the self-play mechanism employed in our method.
2.1. LLMs for Recommendation
LLMs have demonstrated exceptional generative, generalization, and reasoning capabilities in natural language processing, prompting researchers in the recommendation system field to explore their application for personalized recommendations. There are several paradigms for leveraging LLMs in recommendation tasks: (1) using them directly as decision makers (Bao et al., 2023b, 2024), (2) assisting the recommendation process by providing additional information (Sheng et al., 2024; Liu et al., 2024; Geng et al., 2024), and (3) acting as user simulators (Zhang et al., 2024b; Cai et al., 2024; Zhang et al., 2024a). Early research focused on crafting prompts to stimulate the capabilities of LLMs to handle recommendation tasks (Gao et al., 2023b; Hou et al., 2024; Wang et al., 2023b; Gao et al., 2024b).
Subsequently, fine-tuning methods emerged, showing that adapting LLM parameters on recommendation data significantly improves their performance on these tasks. These approaches often rely on Supervised Fine-Tuning (SFT) (Liu et al., 2023; Chen et al., 2025). To further align LLMs with personalized human preferences, researchers adopted Direct Preference Optimization (DPO) for post-training (Chen et al., 2024b; Bai et al., 2024; Liao et al., 2024a). However, no prior work has highlighted the inherent mechanisms of DPO that can introduce severe biases in recommender systems. These biases result in recommendations favoring only the exposed items, ultimately harming user experience. In this work, we are the first to identify this issue and propose a mitigation strategy.
2.2. Biases in Recommender Systems
Bias and unfairness issues are pervasive in recommender systems and have been extensively studied by researchers. Chen et al. (2023a) provide a comprehensive survey summarizing various types of biases, including popularity bias, selection bias, and position bias, among others. When practical recommender systems suffer from these biases, their output can significantly impact user satisfaction. For instance, they may promote content with clickbait titles (Wang et al., 2021) or confine users within filter bubbles, leading to boredom and reduced engagement (Anwar et al., 2024; Areeb et al., 2023; Gao et al., 2023c). Moreover, algorithmic decisions made by recommender platforms can be biased towards certain items, raising concerns about fairness (Wang et al., 2023c; Li et al., 2023). Such biases can disproportionately affect specific user groups or diminish creators’ motivation to produce content (Yao et al., 2024; Jagadeesan et al., 2024).
These issues have persisted throughout the evolution of recommendation systems and continue to exist in the era of LLMs (Wu et al., 2024b; Xu et al., 2024; Tommasel, 2024). Many researchers have identified that LLMs can exhibit social biases, leading to unfair recommendations associated with sensitive attributes such as gender and race (Zhang et al., 2023a; Sakib and Das, 2024b; Shen et al., 2023; Deldjoo and Di Noia, 2024; Ferrara, 2023). Dai et al. (2024) provided a comprehensive review of bias and fairness issues in information retrieval, proposing a unified distribution mismatch perspective to address all forms of bias and unfairness. They also summarized efforts to mitigate bias in LRSs. However, existing efforts either rely on a predefined target distribution or require additional information to guide the alignment direction for LLMs. In this work, we are the first to explore a self-play framework for mitigating bias in LRSs. Unlike previous methods, our approach requires neither prior knowledge nor additional models. By simply altering the tuning procedure, we effectively reduce the long-tail effect and enhance both diversity and fairness in LRSs.
2.3. Self-Play Mechanism
Machine learning models are often data-driven, relying heavily on the availability of offline data. However, offline data is inherently limited, raising an important question: can algorithms improve themselves iteratively without the need for additional data? This is precisely the challenge that the self-play mechanism aims to address. The concept of self-play originated from board games such as Go and chess (Samuel, 1959; Tesauro et al., 1995), exemplified by groundbreaking systems like AlphaGo Zero (Silver et al., 2017b) and AlphaZero (Silver et al., 2017a).
In the era of LLMs, early preference alignment algorithms like RLHF and DPO operate as single optimization procedures. Building on this, Chen et al. (2024a) proposed that LLMs can refine their capabilities through self-play by interacting with instances of themselves. More specifically, the LLM generates its own training data from prior iterations and subsequently learns a new policy that outperforms the old one (Wu et al., 2024a). This iterative fine-tuning framework has been shown to be equivalent to finding the Nash equilibrium of a two-player game (Munos et al., 2023; Rosset et al., 2024; Calandriello et al., 2024), gaining significant recognition due to its solid theoretical foundation and simplicity (Gao et al., 2024a; Ye et al., 2024). In recommender systems, we leverage the self-play mechanism to adaptively reduce biased items in a simple and effective manner.
3. Preliminary
In this section, we provide a brief overview of the technologies for aligning LLMs with the recommendation task. We then introduce the idea of evaluating the biases and unfairness in LRSs.
Figure 2. Illustration of the pipelines and principles of SFT, DPO, and our proposed SPRec in LLM-based recommendations.
3.1. Supervised Fine-tuning (SFT)
To enable an open-sourced LLM to effectively learn recommendation tasks, one of the most practical approaches is fine-tuning all or a subset of its parameters using a set of demonstration data derived from offline recommendation logs. The fine-tuning process aligns the model’s behavior with the recommendation objective by optimizing its parameters on the given dataset. The objective function for this fine-tuning process is defined as maximizing the log-likelihood over the offline training dataset $\mathcal{D}$:
(1) $\pi_{SFT}=\arg\max_{\pi_{\theta}}\mathbb{E}_{(x_{i},y_{i})\sim\mathcal{D}}\log\pi_{\theta}(y_{i}|x_{i}),$
where $(x_{i},y_{i})$ denotes the input-output pairs sampled from the training dataset $\mathcal{D}$. Here, $x_{i}$ represents the context of a recommendation, typically comprising user features and a sequence of historically interacted items, while $y_{i}$ denotes the target item to be recommended. Let $p_{\mathcal{D}}(y|x)$ be the probability of $y$ given $x$ in the dataset $\mathcal{D}$, i.e., the popularity. SFT’s formulation is equal to learning to align the model’s predictions with it by minimizing the forward KL-divergence $\mathbb{D}_{KL}(p_{\mathcal{D}}(y|x),\pi_{\theta}(y|x))$:
$\pi_{SFT}$ $=\arg\min_{\pi_{\theta}}\mathbb{D}_{KL}(p_{\mathcal{D}}(y|x),\pi_{\theta}(y|x))$
(2) $=\arg\min_{\pi_{\theta}}\mathbb{E}_{(x_{i},y_{i})\sim D}-\log\pi_{\theta}(y_{i}|x_{i})+H(p_{\mathcal{D}}),$
where $H(p_{\mathcal{D}})$ denotes the entropy of $p_{\mathcal{D}}$, which is constant with respect to the optimization of the LLM parameters $\theta$.
However, relying exclusively on subsequent targets as positive samples may confuse the model about the true objectives. In recommendation tasks, negative samples play a critical role in highlighting undesirable outcomes (Shi et al., 2023). Thus, it is natural to consider incorporating both positive and negative samples from the offline data to better align the model with user preferences.
3.2. RLHF and DPO
To ensure that outputs align with intricate user preferences (such as helpfulness and harmlessness (Bai et al., 2022)), researchers have introduced Reinforcement Learning with Human Feedback (RLHF) to train LLMs using reward signals (Ouyang et al., 2022). Typically, the RLHF stage follows the SFT stage, further refining language models to better adhere to human instructions. RLHF involves learning a reward model $r_{\phi}(x,y)$, which captures human preferences. The optimization objective during this stage is defined as:
(3) $\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(y|x)}[r_{\phi}(x,y)]-\beta\mathbb{D}_{KL}[\pi_{\theta}(y|x),\pi_{\mathrm{ref}}(y|x)],$
where $\beta$ is a regularization term that penalizes deviations of the learned model $\pi_{\theta}$ from the reference policy $\pi_{\mathrm{ref}}$.
While RLHF has demonstrated remarkable success in aligning LLMs with human preferences, it is computationally intensive, complex, and heavily reliant on the reward model for effectiveness. To simplify reward modeling and mitigate the instability associated with RLHF training, researchers have proposed Direct Preference Optimization (DPO) (Rafailov et al., 2024), which utilizes a closed-form solution of RLHF’s optimal policy. The resulting objective function for DPO is:
(4) $\min_{\pi_{\theta}}-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\log\sigma\Bigg{[}\!\beta\log\!\left(\!\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\mathrm{ref}}(y_{w}|x)}\!\right)-\beta\log\!\left(\!\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\mathrm{ref}}(y_{l}|x)}\!\right)\!\Bigg{]},$
where $(x,y_{w},y_{l})$ represents a prompt $x$ alongside a chosen (preferred) answer $y_{w}$ and a rejected answer $y_{l}$. In the context of recommendation tasks, $x$ corresponds to the recommendation context, as in the SFT stage, while $y_{w}$ and $y_{l}$ denote a positive and a negative sample, respectively.
DPO offers an efficient and stable alternative to RLHF, aligning models with user preferences while reducing reliance on complex reward modeling. Its ability to naturally leverage positive and negative samples makes it particularly well-suited for recommender systems (Chen et al., 2024b; Bai et al., 2024; Liao et al., 2024a).
3.3. Evaluating Biases and Unfairness through Distribution Alignment Perspective
When aligning user preferences, LLMs may inadvertently learn biased or unfair outcomes. To assess the bias and fairness issues in LRSs, a mainstream perspective is to formulate the problems as a mismatch distribution problem (Dai et al., 2024). Specifically, let $R$ denote the ground-truth user preference (e.g., an item list), following the distribution $P(R)$, and let $\hat{R}$ represent the model-predicted preferences, drawn from $P(\hat{R})$. Bias or unfairness is then quantified by the mismatch between these two distributions:
(5) $P(R)\neq P(\hat{R}).$
To operationalize this framework, we adopt the method outlined by Jiang et al. (2024), where $P(R)$, the target user history preference distribution, is approximated using the category distribution derived from offline training data. Furthermore, we also employ their proposed metrics, MGU@$K$ and DGU@$K$, to systematically measure the degree of mismatch in our experiments.
4. Method
In this section, we first point out the severe popularity bias due to the nature of DPO, then we analyze how we can address this by utilizing the self-play philosophy. Then we detail the proposed SPRec architecture.
4.1. Problem of DPO: Amplify Popular Bias
As we mentioned DPO is effective and efficient in aligning LLMs with the user preferences, it has a problem of suppressing the negative samples too much, which will result in severe popularity bias in the recommendation task. In recommendation tasks, the input and positive samples $(x,y_{w})$ are derived from logged interactions in offline data, while negative samples $y_{l}$ are drawn from non-interacted items. Given an input $x$, the probability $p_{\mathcal{D}}(y|x)$ represents the popularity of item $y$ in the dataset, and $q_{\mathcal{D}}(y|x)$ denotes the probability of the same item $y$ being selected as a negative sample. Using these definitions, the DPO loss in Eq. (4) can be rewritten as:
(6) $\begin{gathered}\mathcal{L}_{\mathrm{DPO}}(\pi_{\theta};\pi_{ref})=-\mathbb{E}_{(x,y_{w})\sim\mathcal{D},y_{l}\sim q_{\mathcal{D}}(\cdot|x)}\ell(\pi_{\theta},\pi_{\mathrm{ref}},x,y_{w},y_{l}),\\ \text{with }\ell(\cdot)=\log\sigma[\beta\log(\frac{\pi_{\theta}(y_{w}|x)}{\pi_{ref}(y_{w}|x)})-\beta\log(\frac{\pi_{\theta}(y_{l}|x)}{\pi_{ref}(y_{l}|x)})].\end{gathered}$
In this setting, the optimal policy has a closed-form solution, as stated in the following theorem:
Theorem 1.
The optimal policy $\pi_{\theta}^{*}(\cdot|x)$ for the DPO loss defined in Eq. (6) is given by:
$\pi_{\theta}^{*}(y|x)\propto\pi_{\mathrm{ref}}(y|x)\cdot\left(\frac{p_{\mathcal{D}}(y|x)}{q_{\mathcal{D}}(y|x)}\right)^{1/\beta}.$
The proof is deferred to Appendix A. This result highlights that the optimal policy is proportional to the reference policy $\pi_{\mathrm{ref}}(y|x)$, adjusted by the relative likelihood ratio $\left(\frac{p_{\mathcal{D}}(y|x)}{q_{\mathcal{D}}(y|x)}\right)^{1/\beta}$.
In most recommendation setting, the probability of an item $y$ being a negative sample satisfies a uniform distribution (Bai et al., 2024; Chen et al., 2024b), i.e., $q_{\mathcal{D}}(y|x)=\mathcal{U}=\frac{1}{|\mathcal{I}|}$, where $\mathcal{I}$ is the all item candidates. Furthermore, in most preference alignment scenarios using the DPO-based method, the parameter $\beta$ is constrained to the range $10<\beta<$ 1. Under these conditions, the DPO loss naturally encourages the policy to prioritize popular items with higher probabilities $p_{\mathcal{D}}(y|x)$, thereby introducing a notable popularity bias.
In the extreme case where $\beta\to 0$, the optimal policy under the DPO loss degenerates to recommending only the most popular items with the highest probabilities $p_{\mathcal{D}}(y|x)$, effectively excluding less popular but potentially valuable items from consideration.
Remark: This result is a byproduct of DPO’s loss function. Unlike the forward KL-divergence $\mathbb{D}_{KL}(p_{\mathcal{D}}(y|x),\pi_{\theta}(y|x))$ used in the SFT loss in Eq. (3.1), DPO optimizes the reverse KL-divergence $\mathbb{D}_{KL}(\pi_{\theta}(y|x),\pi_{\mathrm{ref}}(y|x))$. Forward KL-divergence is known for its mass-covering property, which encourages learning an average behavior and is less sensitive to subtle differences in the preference distribution (as illustrated in Fig. 2(a)) (Sun and van der Schaar, 2024). In contrast, the reverse KL-divergence used in DPO promotes mode-seeking behavior (Wang et al., 2023a; Omura et al., 2024), guiding the model to focus on the “peaks” of the distribution (Fig. 2(b)).
The issue of DPO has also been highlighted by other researchers (Pal et al., 2024). Specifically, Feng et al. (2024) derive that DPO suppresses negative samples more aggressively than it elevates positive samples during optimization. Moreover, Azar et al. (2024) demonstrate that the empirical optimal policy often drives $\pi_{\theta}(y_{l}|x)\rightarrow 0$ for all $\beta$, stemming from an underfitting of the potential reward.
In the context of recommendation, this behavior can be detrimental, as it exacerbates the filter bubble issue and undermines user interests by limiting exposure to diverse items (Gao et al., 2023a).
4.2. Solution: Suppress Biases through Self-Play
Since the DPO loss inherently causes the policy $\pi_{\theta}$ to learn sharp ”peaks”, leading to bias, an intuitive solution is to directly suppress these learned peaks. To address this, we utilize a self-play framework, dubbed SPDPO, which iteratively alternates between policy learning and bias suppression. Specifically, in the $(t+1)$-th iteration, negative samples are drawn from the model’s predictive distribution $\pi_{\theta_{t}}(\cdot|x)$ at iteration $t$, resulting in the following learning paradigm:
(7) $\pi_{\theta_{t+1}}\leftarrow\arg\max_{\pi_{\theta}}\mathbb{E}_{(x,y_{w})\sim D,y_{l}\sim\pi_{\theta_{t}}(\cdot|x)}l(\pi_{\theta};\pi_{\theta_{t}};x,y_{w},y_{l}).$
By comparing Eq. (7) with Eq. (6), we obtain that the objective function $\mathcal{L}_{SPDPO}$ in the $(t+1)$-th iteration can be viewed as $\mathcal{L}_{DPO}$ weighted by $\frac{\pi_{\theta_{t}}(y_{l}|x)}{q_{\mathcal{D}}(y_{l}|x)}$, which can be expressed as follows:
(8) $\mathcal{L}_{SPDPO}=-\mathbb{E}_{(x,y_{w})\sim\mathcal{D},y_{l}\sim q_{\mathcal{D}}(\cdot|x)}\frac{\pi_{\theta_{t}}(y_{l}|x)}{q_{\mathcal{D}}(y_{l}|x)}l(\pi_{\theta};\pi_{\theta_{t}};x,y_{w},y_{l}).$
Again, if DPO uses negative samples from a discrete uniform distribution $q_{\mathcal{D}}(y|x)=\mathcal{U}=\frac{1}{|\mathcal{I}|}$, then the objective function $\mathcal{L}_{SPDPO}$ can be viewed as $\mathcal{L}_{DPO}$ weighted by $\pi_{\theta_{t}}(y_{l}|x)$. This highlights that the objective adaptively pays more attention to biased items by increasing their learning rates if they have higher probabilities in the model’s output distribution.
Remark: Unlike traditional recommendation methods that predefine negative samples or allocate weights in advance, our approach dynamically selects negative samples during the learning process. This provides a significant advantage, enabling the model to adaptively adjust its learning paradigm for effective bias suppression. As a result, this approach mitigates the filter bubble issue and enhances the diversity of recommendations.
4.3. Architecture of SPRec
Utilizing the loss function in Eq. (7), we propose a self-play recommendation tuning framework, SPRec, which generally includes multiple iterations of both an SFT step and a DPO step. The workflow is illustrated in Fig. 2(c), in which three key steps are conducted sequentially in each iteration:
  1. (1)
    Dataset Construction: For each positive sample $\{(x^{i},y_{w}^{i})\}$ in the offline dataset, sample a negative sample $y_{l}^{i}$ by running the current model $\pi_{\theta_{t}}$ and using its predicted recommendation as $y_{l}^{i}$. Thus we obtain pairwise preference data for each sample as $\{(x^{i},y_{w}^{i},y_{l}^{i})\}$.
  2. (2)
    SFT Step: Use only the positive sample $\{(x^{i},y_{w}^{i})\}$ to refine the model $\pi_{\theta_{t}}$ through SFT techniques such as instruction learning.
  3. (3)
    DPO Step: Align the model $\pi_{\theta_{t}}$ by perform DPO step using pairwise dataset $\{(x^{i},y_{w}^{i},y_{l}^{i})\}$, and obtain $\pi_{\theta_{t+1}}$.
The iteration will repeat $T$ times in an epoch. The self-play philosophy can utilized in any LLM tuning-based LRSs. Next, we will explain the necessity of the SFT step, which is not considered in original self-play fine-tuning work (Chen et al., 2024a).
Remark: Although the loss function in Eq. (7) is inherently capable of aligning the model with positive samples, our experiments reveal that incorporating an SFT step in each self-play iteration can further enhance performance.
In fact, combining SFT and DPO has been shown to be an effective practice in recent research and open-sourced LLM models. For example, each iteration of post-training for Llama 3 includes an SFT stage followed by a DPO stage (Dubey et al., 2024). Similarly, Pang et al. (2024) demonstrate the effectiveness of an iterative preference optimization algorithm useing a modified DPO loss with an additional negative log-likelihood (NLL) term, which mirrors the SFT loss defined in Eq. (1).
Table 1. Statistics of Datasets.
Datasets #Items #Interactions #Sequences
MovieLens 10,682 10,000,054 9,301,274
Goodreads 4,058 160,398 6,031
CDs_and_Vinyl 13,078 185,855 21,347
Steam 32,094 178,961 29,876
5. Experiments
In this section, we conduct experiments to address the following research questions:
  • RQ1: How does the SPRec training framework compare to baseline methods in terms of accuracy, diversity, and fairness?
  • RQ2: What are the contributions of different components within the SPRec framework?
  • RQ3: How do the random sampling ratio and the number of negative samples impact the performance?
Table 2. Overall performance comparison of SPRec (green) with SFT-based (brown) and DPO-based (blue) methods. Best results are in bold, sub-optimal results are underlined. Metrics with $\uparrow$ indicate higher is better, while $\downarrow$ indicates lower is better.
Dataset Model DivRatio@1$\uparrow$ DivRatio@5$\uparrow$ ORRatio@1$\downarrow$ ORRatio@5$\downarrow$ MGU@1$\downarrow$ MGU@5$\downarrow$ HR@5$\uparrow$ NDCG@5$\uparrow$
SASRec 0.0133 0.0057 0.6588 0.4778 0.0463 0.0439 0.0240 0.0155
[…]
RW 0.1918 0.1541 0.2551 0.0990 0.0577 0.0405 0.0327 0.0276
$\text{D}^{3}$ 0.1246 0.1000 0.3238 0.1440 0.0664 0.0502 0.0266 0.0197
[…]
SASRec 0.3236 0.1889 0.1391 0.1149 0.0070 0.0063 0.1149 0.0837
[…]
RW 0.1050 0.0820 0.3730 0.1338 0.0238 0.0270 0.0380 0.0281
[…]
SASRec 0.0021 0.0035 1.0000 0.5529 0.1446 0.0632 0.0230 0.0161
[…]
RW 0.2821 0.2646 0.2648 0.0971 0.0432 0.0305 0.0112 0.0103
$\text{D}^{3}$ 0.3135 0.2968 0.2672 0.0748 0.0320 0.0252 0.0103 0.0088
[…]
SASRec 0.3640 0.1794 0.1610 0.1102 0.0241 0.0209 0.1940 0.1407
[…]
$\text{D}^{3}$ 0.1580 0.1414 0.4560 0.1896 0.0511 0.0320 0.0730 0.0718
[…]
SDPO 0.2140 0.1978 0.3270 0.0756 0.0448 0.0320 0.0810 0.0804
RosePO 0.2060 0.1868 0.3670 0.1080 0.0562 0.0349 0.0760 0.0750
[…]
5.1. Experimental Setup
5.1.1. Datasets
We conducted extensive experiments on four real-world datasets: MovieLens [^1] https://grouplens.org/datasets/movielens/ (Harper and Konstan, 2015), Steam [^2] https://cseweb.ucsd.edu/j̃mcauley/datasets.html#amazon_reviews, Goodreads [^3] https://mengtingwan.github.io/data/goodreads, and the CDs and Vinyl category of the Amazon Review Dataset [^4] http://jmcauley.ucsd.edu/data/amazon/index_2014.html. Following the data processing approach in (Chen et al., 2024b; Bao et al., 2024), interaction sequences with fewer than 10 entries were excluded. The datasets were then split chronologically into training, validation, and test sets in an 8:1:1 ratio, ensuring mutual exclusivity and preventing data leakage. The statistics of the datasets are summarized in Table 1.
To process category information, we extracted category metadata from each dataset and identified the most $10$ popular categories within the training sets. To ensure category independence, we removed categories with clear hierarchical relationships, such as “FPS” and “Shooting” in the Steam dataset, and “Rock” and “Classical Rock” in the CDs and Vinyl dataset.
Table 3. Ablation Experiment Results. SPRec-RN represents the training process with negative samples randomly sampled, while SPRec-noDPO and SPRec-noSFT denote SPRec without DPO and SFT, respectively.
Dataset Model DivRatio@1$\uparrow$ DivRatio@5$\uparrow$ ORRatio@1$\downarrow$ ORRatio@5$\downarrow$ MGU@1$\downarrow$ MGU@5$\downarrow$ HR@5$\uparrow$ NDCG@5$\uparrow$
SPRec-noSFT 0.3020 0.2486 0.0837 0.0224 0.0198 0.0188 0.0184 0.0149
[…]
SPRec-noDPO 0.2220 0.2006 0.3910 0.1884 0.0620 0.0294 0.0790 0.0776
[…]
Figure 3. Comparison of models across genres on Group Unfairness (GU) in top-$1$ recommendation.
[…]
Fig: The Group Unfairness (GU) of different groups divided by genres in top-1 recommendation.
5.1.2. Evaluation Setting
To leverage the strengths of LLMs in generative recommendation tasks, we prompt the LLM to generate a predicted item based on the input history sequence. Then, following the procedures in BIGRec (Bao et al., 2023a), we calculate scores and rankings for the entire item space and ground our predicted item to an exact item in the dataset.
5.1.3. Metrics.
We evaluate the model using four key metrics on a randomly sampled set of 1,000 test cases per iteration. Accuracy is measured using Normalized Discounted Cumulative Gain (NDCG@$k$) and Hit Ratio (HR@$k$) and calculated per result before averaging. Diversity is assessed with DivRatio@$k$, representing the proportion of unique items in the top-$k$ recommendations. Over-recommendation is quantified by ORRatio@$k$, indicating the proportion of the top-$k$ results consisting of the three most frequently recommended items. Fairness is evaluated using MGU@$k$ (Jiang et al., 2024), which captures category-level discrepancies between the model’s top-$k$ recommendations and the user’s historical sequences.
5.1.4. Baseline
For traditional recommendation models, we select SASRec (Kang and McAuley, 2018), a widely used baseline employing a sequential method with a self-attention mechanism. For LLM-based models, we consider several baselines. (1) For SFT-based methods, BIGRec (Bao et al., 2023a) serves as an instruction-tuning LLM framework for sequential recommendations and forms the foundation for SPRec. Re-weighting (RW) (Jiang et al., 2024) improves fairness in BIGRec by balancing recommendations across categories through dataset-based training weights. Debiasing-Diversifying Decoding ($\textbf{D}^{3}$) (Bao et al., 2024) enhances diversity in BIGRec using a decoding strategy guided by SASRec. (2) For DPO-based models, DMPO (Bai et al., 2024) introduces DPO into LRSs by sampling multiple negative items as rejected responses, while Softmax-DPO (SDPO) (Chen et al., 2024b) follows a similar approach but incorporates a softmax loss over multiple negative samples. Finally, RosePO (Liao et al., 2024a) is a preference optimization framework that combines negative sampling strategies and personalized uncertainty to achieve fairness, unbiasedness, and robustness. The implementation details are listed in Appendix B.
5.2. Overall Performance Comparison (RQ1)
The experimental results are shown in Table 5, leading to the following observations.
5.2.1. Limitations of SFT-based Methods
Fine-tuning LLMs with instruction-based methods results in recommendations heavily favoring popular items, leading to a lack of diversity. For example, in the Goodreads dataset, the DivRatio@1 of BIGRec is only 0.142, meaning the model provides just 14 distinct recommendations per 100 tasks. Similarly, in the Steam dataset, BIGRec’s ORRatio@1 reaches 0.391, with over 39% of recommendations concentrated on the 3 most popular items. These findings highlight that relying solely on SFT introduces severe biases, significantly overexposing certain popular items.
5.2.2. Limitations of DPO-based Methods
For DPO methods using random sampling, such as SDPO and DMPO, while multiple negative samples improve recommendation accuracy, they perform poorly on diversity and fairness metrics. In the Goodreads and MovieLens datasets, SDPO and DMPO exhibit significantly higher ORRatio values than BIGRec, supporting the conclusion in Section 4.1 that DPO amplifies popularity bias. On the CDs and Steam datasets, although ORRatio improves, diversity metrics remain largely unchanged, indicating that the model favors moderately popular items but fails to effectively recommend new ones. In summary, existing DPO-based methods not only fail to address fairness issues in LRS but may exacerbate popularity bias.
5.2.3. Superiority of SPRec
As shown in Table 5, SPRec significantly improves both DivRatio and ORRatio metrics across all datasets compared to BIGRec, demonstrating its effectiveness in mitigating the over-recommendation of popular items and enhancing diversity. Additionally, SPRec outperforms BIGRec on most fairness metrics, reducing the discrepancies between the model’s recommendations and users’ historical sequences, thereby providing fairer recommendations.
SPRec also surpasses most baseline models on DivRatio and ORRatio across the majority of datasets, showcasing its superior ability to balance recommendation distributions. For fairness, SPRec achieves the best MGU@1 on all datasets except Steam. Moreover, as shown in Fig. 3, SPRec alleviates category-level unfairness on the MovieLens dataset, achieving the best results in 7 out of 8 categories, further underscoring its effectiveness in improving fairness.
While RosePO performs well on the CDs and Vinyl dataset, leveraging semantic-based negative sampling to address fairness in music recommendations, and Re-weighting shows strong performance on the Steam dataset by employing category-based re-weighting for gaming recommendations, these methods are tailored to specific datasets and lack generalizability. In contrast, SPRec’s self-play framework provides a universal solution, overcoming dataset-specific challenges and delivering fairer recommendations across diverse scenarios.
Figure 4. Performance on the MovieLens dataset across different ablation experiments.
5.3. Ablation Study (RQ2)
We conducted a series of ablation experiments to explore the impact of each component of the SPRec training framework.
5.3.1. SPRec without SFT
As shown in Table 3, SPRec-noSFT achieves the lowest recommendation accuracy across all datasets except Steam. This indicates that, during the self-play process, the model’s excessive focus on fairness compromises its accuracy. In the MovieLens dataset (Fig. 4), the absence of the SFT stage leads to a steady decline in recommendation accuracy (NDCG@1) throughout training. These findings highlight the critical role of SFT in maintaining SPRec’s recommendation performance.
5.3.2. SPRec without DPO
Removing DPO reduces SPRec to further SFT training on new data, ensuring that performance gains are not due to incorporating additional data. As shown in Table 3, even with new data, additional SFT fails to improve diversity or fairness metrics, and the recommendations remain biased. Furthermore, Fig. 4 reveals minimal fluctuations during training, indicating that the prior SFT training has already converged. This experiment underscores the limitations of SFT-based methods in addressing recommendation fairness and diversity.
5.3.3. Randomly sampling negative items
Replacing the negative sampling strategy with random sampling (SPRec-RN) fails to improve DivRatio and ORRatio metrics on the MovieLens and Goodreads datasets (Table 3). Additionally, SPRec-RN exhibits worse fairness metrics compared to SPRec. While it achieves a significant boost in DivRatio on the CDs dataset, its ORRatio remains poor. These results suggest that random negative sampling is ineffective in suppressing popular items, leaving a pronounced long-tail effect in the recommendations. This experiment demonstrates that our self-play negative sampling strategy effectively balances the model’s output distribution, resulting in debiased recommendations.
Figure 5. Effect of random sampling ratio.
Figure 6. Effect of negative sample size.
5.4. Impact of Negative Samples (RQ3)
We investigate the role of negative samples in SPRec by introducing a proportion of random negative samples to contaminate SPRec’s original self-play samples. Additionally, we examine the impact of increasing the number of negative samples in SPRec’s loss function (Eq. (8)). To achieve this, we adopt the SDPO loss function to expand to multiple negative samples. For efficiency, we limit the training sample size to 1,024, keeping other experimental settings consistent with Section 5.2. To ensure result stability, the model’s performance is averaged over the last three training iterations.
5.4.1. Effect of Random Sampling Ratio
In each training iteration, we randomly replace a proportion of negative samples with randomly selected items while leaving the remaining negative samples unchanged as the model’s recommendation outputs. As shown in Fig. 5, increasing the proportion of random negative samples leads to a steady decline in diversity and accuracy. Fairness also deteriorates, with recommendations becoming more skewed toward a small subset of popular items. These results highlight the superiority of our proposed self-play negative sampling strategy over random sampling.
5.4.2. Effect of Negative Sample Size
To generate $N$ negative samples, we use beam search decoding to sample $2N$ items from the model’s output. After deduplication, the top $N$ items with the highest predicted probabilities are selected as negative samples. As shown in Fig. 6, increasing the number of negative samples results in stable recommendation accuracy but significantly improves diversity and fairness, reducing the focus on popular items. This demonstrates the versatility of the self-play negative sampling strategy, which can be effectively combined with multi-negative sampling approaches to further debias LRS.
6. Conclusion
This paper theoretically and experimentally analyzes the limitations of existing SFT- and DPO-based methods in LLM-based recommender systems. These approaches often promote a narrow set of items, intensifying the filter bubble effect, harming user interests, and raising fairness concerns that may undermine creators’ motivation. To address these challenges, we propose SPRec, a novel self-play recommendation framework that mitigates over-recommendation and enhances fairness without requiring additional data or manual intervention. SPRec leverages the model’s output from the previous round as negative samples, effectively re-weighting the DPO loss function with the model’s logits. This adaptive mechanism suppresses biased items and improves recommendation quality.
We conducted experiments on four real-world datasets, validating the effectiveness of our method in improving recommendation accuracy, diversity, and fairness. Comprehensive ablation studies revealed that fairness improvements largely stem from the suppression of self-play negative samples during DPO. In the future, we plan to explore more advanced preference alignment techniques for LRSs, aiming for smarter personalized decision-making and more helpful and harmless recommendation systems.
Appendix A Mathematical Derivations
Proof of Theorem 1.
The DPO loss is derived from the RLHF objective as follows:
(9) $\max_{\theta}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(\cdot|x)}[r(x,y)]-\beta\mathrm{D_{KL}}(\pi_{\theta}\|\pi_{\mathrm{ref}}),$
where the reward model is defined via the BT model: max_r E_(x,y_w)∼D, y_l∼Y_ulogσ(r(x,y_w) - r(x,y_l)).
In the original paper of DPO (Rafailov et al., 2024), the authors proved that the optimal policy $\pi_{\theta}^{*}$ for DPO loss in Eq. (LABEL:eqn:dpo_def) and the solution to the optimization problem in Eq. (9) are the same. Thus, we can analyze the solution to Eq. (9), equivalent to examining the DPO loss.
Consider a fixed context $x$ and define $\hat{r}(y|x)$ as:
(10) $\hat{r}(y|x)=\exp(r(x,y)),$
then our goal is to optimize ${\boldsymbol{\hat{r}}(\cdot|x)\in\mathbb{R}^{\mathcal{I}}_{+}}$, which is a $|\mathcal{I}|$-dim vector representing the latent rewards for all items in the recommendation dataset $\mathcal{I}$. we can rewrite the reward model’s optimization as: max_^r(⋅—x)∈R^I_+ ∑_y_w∈I∑_y_l∈Ip_D(y_w—x)q_D(y_l—x)log(^r(yw—x)^r(yw—x)+^r(yl—x)). Then we calculate the gradients:
$\partial p_{\mathcal{D}}(y_{w}|x)q_{\mathcal{D}}(y_{l}|x)\log\left(\frac{\hat{r}(y_{w}|x)}{\hat{r}(y_{w}|x)+\hat{r}(y_{l}|x)}\right)/\partial\hat{r}(y|x)$
[…]
Hence, the objective’s gradient w.r.t. $\hat{r}(y|x)$ can be written as:
$\partial\sum_{y_{w}\in\mathcal{I}}\sum_{y_{l}\in\mathcal{I}}p_{\mathcal{D}}(y_{w}|x)q_{\mathcal{D}}(y_{l}|x)\log\left(\frac{\hat{r}(y_{w}|x)}{\hat{r}(y_{w}|x)+\hat{r}(y_{l}|x)}\right)/\partial\hat{r}(y|x)$
$=$ $0+\sum_{y_{l}\neq y}\left[p_{\mathcal{D}}(y|x)q_{\mathcal{D}}(y_{l}|x)(\frac{1}{\hat{r}(y|x)}-\frac{1}{\hat{r}(y|x)+\hat{r}(y_{l}|x)})\right]$
$\phantom{0+}-\sum_{y_{w}\neq y}p_{\mathcal{D}}(y_{w}|x)q_{\mathcal{D}}(y|x)\frac{1}{\hat{r}(y_{w}|x)+\hat{r}(y|x)}$
$=$ $\sum_{y_{i}\in\mathcal{I}}\left[p_{\mathcal{D}}(y|x)q_{\mathcal{D}}(y_{i}|x)(\frac{1}{\hat{r}(y|x)}-\frac{1}{\hat{r}(y|x)+\hat{r}(y_{i}|x)})\right]$
[…]
$=$ $\sum_{y_{i}\in\mathcal{I}}\left[\frac{p_{\mathcal{D}}(y|x)q_{\mathcal{D}}(y_{i}|x)}{\hat{r}(y|x)}-\frac{p_{\mathcal{D}}(y_{i}|x)q_{\mathcal{D}}(y|x)+p_{\mathcal{D}}(y|x)q_{\mathcal{D}}(y_{i}|x)}{\hat{r}(y_{i}|x)+\hat{r}(y|x)}\right].$
By setting the gradients to be 00, we obtain that for $\forall y\in\mathcal{I}$, the optimal reward $r^{*}(y|x)$ is: ^r^*(y—x) ∝pD(y—x)qD(y—x).
By plugging it into Eq. (10), we have: r^*(x,y) = logp_D(y—x) - logq_D(y—x) + Constant.
Back to the RLHF objective, we have:
$\max_{\theta}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(\cdot|x)}[\log p_{\mathcal{D}}(y|x)-\log q_{\mathcal{D}}(y|x)]-\beta\mathrm{D_{KL}}(\pi_{\theta}\|\pi_{\mathrm{ref}}),$
which has a well-known closed form solution (Rafailov et al., 2024): π_θ^*(y—x) ∝π_ref(y—x)⋅(pD(y—x)qD(y—x))^1/β. ∎
Appendix B Implementation Details
For all traditional models, we set the sizes of the training and validation sets to $65,536$ and $8,192$, respectively. The embedding size was fixed at $64$, and the dropout ratio was set to 0.1. Negative samples were randomly sampled in training, with Adam as the optimizer and a learning rate of 4e-3. For LLM-based methods, we adopted Llama-3.2-1B-Instruct as the backbone LLM. Considering the ability of LLMs to quickly adapt to downstream tasks with limited data, we followed BIGRec (Bao et al., 2023a) and used relatively smaller datasets compared to traditional models. For SFT-based methods, including BIGRec and Re-weighting, the sizes of the training and validation sets were set to $4,096$ and $512$, respectively. For DPO-based methods, the sizes were set to $2,048$ and $256$, and training was uniformly conducted on the basis of BIGRec. For SPRec, the total number of iterations was set to $5$, with each SFT and DPO phase trained for 1 epoch. All experiments were carried out on four RTX 3090 GPUs, each with 24GB of VRAM.

### v3

@@
-SPRec: Leveraging Self-Play to Debias Preference Alignment for Large Language Model-based Recommendations
+SPRec: Self-Play to Debias LLM-based Recommendation
 Abstract.
 Large language models (LLMs) have attracted significant attention in recommendation systems.
-Current LLM-based recommender systems primarily rely on supervised fine-tuning (SFT) to train the model for recommendation tasks.
-However, relying solely on positive samples limits the model’s ability to align with user satisfaction and expectations.
-To address this, researchers have introduced Direct Preference Optimization (DPO), which explicitly aligns recommendations with user preferences using offline preference ranking data.
-Despite its advantages, our theoretical analysis reveals that DPO inherently biases the model towards a few items, exacerbating the filter bubble issue and ultimately degrading user experience.
-In this paper, we propose SPRec, a novel self-play recommendation framework designed to mitigate over-recommendation and improve fairness without requiring additional data or manual intervention.
+Current work primarily applies supervised fine-tuning (SFT) to adapt the model for recommendation tasks.
+However, SFT on positive examples only limits the model’s ability to align with user preference.
+To address this, researchers recently introduced Direct Preference Optimization (DPO), which explicitly aligns LLMs with user preferences using offline preference ranking data.
+However, we found that DPO inherently biases the model towards a few items, exacerbating the filter bubble issue and ultimately degrading user experience.
+In this paper, we propose SPRec, a novel self-play framework designed to mitigate over-recommendation and improve fairness without requiring additional data or manual intervention.
 In each self-play iteration, the model undergoes an SFT step followed by a DPO step, treating offline interaction data as positive samples and the predicted outputs from the previous iteration as negative samples.
 This effectively re-weights the DPO loss function using the model’s logits, adaptively suppressing biased items.
-Extensive experiments on multiple real-world datasets demonstrate SPRec’s effectiveness in enhancing recommendation accuracy and addressing fairness concerns.
+Extensive experiments on multiple real-world datasets demonstrate SPRec’s effectiveness in enhancing recommendation accuracy and fairness.
+The code is available via https://github.com/RegionCh/SPRec.
 Large Language Model-based Recommendation, Homogeneity Issue, Fairness, Self-play, Direct Preference Optimization
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
+copyright: acmlicensedjournalyear: 2025copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714524isbn: 979-8-4007-1274-6/25/04ccs: Information systems Recommender systems
 1. Introduction
-Recently, large language models (LLMs) have demonstrated significant success across numerous domains, showcasing advanced capabilities in learning, reasoning, and generalizing to downstream tasks (Zhao et al., 2023; Wang et al., 2024a; Dai et al., 2024).
-In the field of recommender systems, there has been growing interest in leveraging the potential of LLMs in novel ways (Wu et al., 2024c).
-One prominent approach involves positioning LLMs as the central recommendation backbone, utilizing users’ past interactions and current needs to generate personalized item recommendations (Bao et al., 2023a; Wei et al., 2024).
-Compared to traditional methods, LLM-based recommendation systems (LRSs) offer distinct advantages, including a deeper contextual understanding and the flexibility to dynamically adapt to users’ evolving preferences.
+Recently, large language models (LLMs) have demonstrated significant success across numerous domains, showcasing advanced capabilities in learning, reasoning, and generalizing to downstream tasks (Wang et al., 2024b; Dai et al., 2024).
+In the field of recommender systems, there has been growing interest in leveraging the potential of LLMs (Wu et al., 2024).
+One prominent approach involves positioning LLMs as the central recommendation backbone, utilizing users’ past interactions and current needs to generate personalized recommendations (Bao et al., 2025; Wei et al., 2024).
+Compared to traditional methods, LLM-based recommendation systems (LRSs) offer distinct advantages, including a deeper contextual understanding and the flexibility to adapt to users’ evolving preferences.
 Figure 1. Homogeneity issues in LLM-based recommendation results caused by token-level and item-level biases.
-To enable LLMs to learn collaborative filtering signals and effectively perform specific recommendation tasks, a prevalent strategy is to fine-tune pre-trained LLMs via Supervised Fine-Tuning (SFT) on offline recommendation datasets (Bao et al., 2023a, b; Liao et al., 2024b; Zhang et al., 2023b).
-This approach allows LLMs to efficiently internalize user preferences within the offline data by adjusting their parameters to align with the recommendation task.
-Building on the foundation of SFT, recent research has adopted Direct Preference Optimization (DPO) (Rafailov et al., 2024) to further refine user preferences (Chen et al., 2024b; Bai et al., 2024; Liao et al., 2024a).
-While SFT relies solely on desirable answers, DPO incorporates both chosen and rejected response pairs, allowing the LLM to learn user ranking preferences and develop a more nuanced understanding of fine-grained, personalized information.
-This approach aligns with the common practice in recommendation models, which utilize both positive and negative samples to train the model effectively (Chen et al., 2023b; Shi et al., 2023).
-Despite these advancements, we identify that employing DPO to align user preferences in recommender systems inherently introduces significant biases due to its underlying mechanisms.
-These biases can lead to serious homogeneity issues, where LLMs produce items with similar names or content.
+To enable LLMs to learn collaborative filtering signals and effectively perform item recommendations, a prevalent strategy is to fine-tune pre-trained LLMs via Supervised Fine-Tuning (SFT) (Bao et al., 2025).
+This approach allows LLMs to efficiently internalize user preferences from offline data by adjusting their parameters to align with the recommendation task.
+Building on SFT, recent research has adopted Direct Preference Optimization (DPO) (Rafailov et al., 2024) to further refine user preferences (Chen et al., 2024b; Bai et al., 2024; Liao et al., 2024).
+While SFT relies solely on desirable answers, DPO incorporates both chosen and rejected response pairs, allowing the LLM to learn user ranking preferences and gain a more nuanced understanding of fine-grained, personalized information.
+This approach mirrors the common practice in recommendation models, which utilize both positive and negative samples for effective training (Chen et al., 2023b; Shi et al., 2023).
+Despite these advancements, we find that employing DPO to align user preferences in recommender systems inherently introduces significant biases due to its underlying mechanisms.
+These biases can lead to serious homogeneity issues, where LLMs recommend items with similar names or content.
 Fig.
-1 provides a simplified visualization of how token-level and item-level biases manifest in the Top-$K$ movie recommendation results.
-Token-level biases occur due to the fact that LLMs generate item names in a tokenized fashion.
-Since LLMs are usually tuned to maximize the likelihood of target tokens, it can result in an overemphasis on items composed of more common tokens.
-For instance, movies with common words like “the” in their titles may be disproportionately recommended, regardless of their actual relevance to user preferences.
+1 illustrates how token-level and item-level biases manifest in the Top-$K$ movie recommendation.
+Token-level biases arise due to the fact that LLMs generate item names in a tokenized fashion.
+Since LLMs are usually tuned to maximize the likelihood of target tokens, items with more common tokens (e.g., movies with the word “the” in their titles) may be overrepresented, regardless of user relevance.
 At the item level, biases can emerge from multiple factors, particularly after fine-tuning, where LLMs may disproportionately recommend popular items, such as the Batman film series.
-This phenomenon can lead to the creation of filter bubbles (Gao et al., 2023a; Wang et al., 2022), where users are repeatedly exposed to a narrow range of highly popular content, ultimately detracting from the diversity of their recommendations and the overall user experience.
-Some research has been conducted on bias and unfairness issues in LRSs (Dai et al., 2024; Zhang et al., 2023a; Wang et al., 2024b; Gallegos et al., 2024; Sakib and Das, 2024a).
+This can lead to filter bubbles (Gao et al., 2023a, c), where users are repeatedly exposed to a narrow range of popular content, limiting the diversity of recommendations and degrading the user experience.
+Some research has been conducted on bias and unfairness issues in LRSs (Dai et al., 2024; Zhang et al., 2023; Gallegos et al., 2024).
 Dai et al.
@@
 For example, Jiang et al.
-(2024) proposed re-weighting the fine-tuning loss for each item and re-ranking the generated results to ensure equal treatment of genre groups.
+(2024) proposed re-weighting the fine-tuning loss for each item and re-ranking the generated results to ensure equitable treatment across genre groups.
 Similarly, Bao et al.
 (2024) adjusted the LLM decoding process by removing the length normalization term and incorporating predictions from a text-free model, helping to reduce amplification bias and address homogeneity issues.
-However, these existing methods either require carefully crafted rules or rely heavily on external knowledge, limiting their applicability in general recommendation systems.
-To this end, we propose a self-play recommendation tuning framework, SPRec, to adaptively suppress biases and improve fairness in LRSs without the need for additional information or expert knowledge.
-The core idea of SPRec is very straightforward: each tuning iteration begins with an SFT round using positive samples from offline data, followed by a DPO round still treating SFT data as positive samples and the predicted outputs from the previous iteration as negative samples.
-The philosophy is to let the model “play” with its own output by re-weighting the DPO loss function based on the predictions.
-Hence, the higher an item ranks in the model’s predictions, the more it is penalized.
-Meanwhile, the SFT process reinforces the ranking of positive items.
-Over time, this self-play learning process adaptively suppresses all undesirable items (biases) while maintaining alignment with positive samples.
-Extensive experiments on public datasets demonstrate that SPRec effectively enhances both accuracy and fairness metrics, showcasing its potential as a practical and efficient solution for LRSs.
+However, these methods often rely on carefully crafted rules or external knowledge, limiting their broader applicability in general recommendation systems.
+To this end, we propose a self-play recommendation tuning framework, SPRec, to adaptively suppress biases and improve fairness in LRSs without the need for additional data or expert knowledge.
+The core idea of SPRec is straightforward: each tuning iteration begins with an SFT round using positive samples from offline data, followed by a DPO.
+In the DPO step, the SFT data is treated as positive samples, while the predicted outputs from the previous iteration are treated as negative samples.
+The philosophy is to let the model “play” with its own output by re-weighting the DPO loss function based on its predictions.
+As a result, items that rank higher in the model’s predictions are penalized, while the SFT process reinforces the ranking of positive items.
+Over time, this self-play learning process adaptively suppresses undesirable items (biases) while maintaining alignment with positive samples.
+Extensive experiments on public datasets demonstrate that SPRec effectively improves both accuracy and fairness, showcasing its potential as a practical and efficient solution for LRSs.
 The main contributions of this paper are as follows:
-  • We provide an analysis that shows current LRSs tuned through DPO inevitably exhibit biases due to their underlying learning mechanisms, leading to the homogeneity issue.
-  • We propose a self-play recommendation tuning framework, SPRec, which addresses these biases and improves fairness without the need for external knowledge.
-  • Experiments validate SPRec’s effectiveness in improving accuracy, diversity, and fairness, with ablation studies attributing the improvements to the self-play negative samples.
+  • We analyze how current LRSs tuned through DPO inevitably exhibit biases due to their underlying learning mechanisms, leading to the homogeneity issue.
+  • We propose SPRec, a self-play recommendation tuning framework that addresses these biases and improves fairness without the need for external knowledge.
+  • Experiments validate SPRec improves accuracy, diversity, and fairness, with ablation studies indicating that the self-play negative samples contribute significantly to the improvements.
 2. Related work
@@
 2.1. LLMs for Recommendation
-LLMs have demonstrated exceptional generative, generalization, and reasoning capabilities in natural language processing, prompting researchers in the recommendation system field to explore their application for personalized recommendations.
-There are several paradigms for leveraging LLMs in recommendation tasks: (1) using them directly as decision makers (Bao et al., 2023b, 2024), (2) assisting the recommendation process by providing additional information (Sheng et al., 2024; Liu et al., 2024; Geng et al., 2024), and (3) acting as user simulators (Zhang et al., 2024b; Cai et al., 2024; Zhang et al., 2024a).
-Early research focused on crafting prompts to stimulate the capabilities of LLMs to handle recommendation tasks (Gao et al., 2023b; Hou et al., 2024; Wang et al., 2023b; Gao et al., 2024b).
-Subsequently, fine-tuning methods emerged, showing that adapting LLM parameters on recommendation data significantly improves their performance on these tasks.
-These approaches often rely on Supervised Fine-Tuning (SFT) (Liu et al., 2023; Chen et al., 2025).
-To further align LLMs with personalized human preferences, researchers adopted Direct Preference Optimization (DPO) for post-training (Chen et al., 2024b; Bai et al., 2024; Liao et al., 2024a).
-However, no prior work has highlighted the inherent mechanisms of DPO that can introduce severe biases in recommender systems.
-These biases result in recommendations favoring only the exposed items, ultimately harming user experience.
+LLMs have shown exceptional generative, generalization, and reasoning capabilities in NLP, driving research into their applications for personalized recommendations.
+Their integration into recommendation tasks follows three main paradigms: (1) acting as decision makers (Bao et al., 2024; Mao et al., 2024), (2) assisting by providing contextual information (Liu et al., 2024; Geng et al., 2024), and (3) serving as user simulators (Zhang et al., 2024; Cai et al., 2024).
+Early studies explored prompt engineering to leverage LLMs for recommendation tasks (Gao et al., 2023b; Hou et al., 2024).
+Later, fine-tuning methods emerged, demonstrating that adapting LLM parameters on recommendation data significantly enhances performance.
+These approaches primarily rely on SFT (Bao et al., 2024; Chen et al., 2025).
+To further align LLMs with user preferences, DPO has been employed for post-training (Chen et al., 2024b; Bai et al., 2024; Liao et al., 2024).
+However, prior work has overlooked DPO’s inherent tendency to introduce severe biases, favoring only frequently exposed items and degrading user experience.
 In this work, we are the first to identify this issue and propose a mitigation strategy.
 2.2. Biases in Recommender Systems
-Bias and unfairness issues are pervasive in recommender systems and have been extensively studied by researchers.
+Bias and fairness issues are pervasive in recommender systems and have been extensively studied.
 Chen et al.
-(2023a) provide a comprehensive survey summarizing various types of biases, including popularity bias, selection bias, and position bias, among others.
-When practical recommender systems suffer from these biases, their output can significantly impact user satisfaction.
-For instance, they may promote content with clickbait titles (Wang et al., 2021) or confine users within filter bubbles, leading to boredom and reduced engagement (Anwar et al., 2024; Areeb et al., 2023; Gao et al., 2023c).
-Moreover, algorithmic decisions made by recommender platforms can be biased towards certain items, raising concerns about fairness (Wang et al., 2023c; Li et al., 2023).
-Such biases can disproportionately affect specific user groups or diminish creators’ motivation to produce content (Yao et al., 2024; Jagadeesan et al., 2024).
-These issues have persisted throughout the evolution of recommendation systems and continue to exist in the era of LLMs (Wu et al., 2024b; Xu et al., 2024; Tommasel, 2024).
-Many researchers have identified that LLMs can exhibit social biases, leading to unfair recommendations associated with sensitive attributes such as gender and race (Zhang et al., 2023a; Sakib and Das, 2024b; Shen et al., 2023; Deldjoo and Di Noia, 2024; Ferrara, 2023).
+(2023a) provide a comprehensive survey on biases such as popularity bias, selection bias, and position bias.
+These biases can significantly impact user satisfaction, promoting clickbait content or reinforcing filter bubbles that reduce engagement (Gao et al., 2023c).
+Additionally, algorithmic decisions may favor certain items, raising fairness concerns (Wang et al., 2023; Li et al., 2023), disproportionately affecting user groups and discouraging content creators (Yao et al., 2024; Jagadeesan et al., 2024).
+These challenges persist in LLM-based recommender systems (Wu et al., 2024; Xu et al., 2024; Tommasel, 2024).
+Research shows that LLMs can inherit social biases, leading to unfair recommendations related to sensitive attributes like gender and race (Zhang et al., 2023).
 Dai et al.
-(2024) provided a comprehensive review of bias and fairness issues in information retrieval, proposing a unified distribution mismatch perspective to address all forms of bias and unfairness.
-They also summarized efforts to mitigate bias in LRSs.
-However, existing efforts either rely on a predefined target distribution or require additional information to guide the alignment direction for LLMs.
-In this work, we are the first to explore a self-play framework for mitigating bias in LRSs.
-Unlike previous methods, our approach requires neither prior knowledge nor additional models.
-By simply altering the tuning procedure, we effectively reduce the long-tail effect and enhance both diversity and fairness in LRSs.
+(2024) provide a unified distribution mismatch perspective on bias and fairness in information retrieval.
+Existing bias mitigation methods in LRSs typically rely on predefined target distributions or external guidance for LLM alignment.
+In contrast, we introduce the first self-play framework for mitigating bias in LRSs, requiring neither prior knowledge nor additional models.
+By simply modifying the tuning process, our approach reduces long-tail effects and improves fairness.
 2.3. Self-Play Mechanism
@@
 This is precisely the challenge that the self-play mechanism aims to address.
-The concept of self-play originated from board games such as Go and chess (Samuel, 1959; Tesauro et al., 1995), exemplified by groundbreaking systems like AlphaGo Zero (Silver et al., 2017b) and AlphaZero (Silver et al., 2017a).
+The concept of self-play originated from board games such as Go and chess, exemplified by groundbreaking systems like AlphaGo Zero (Silver et al., 2017) and AlphaZero (Silver et al., 2018).
 In the era of LLMs, early preference alignment algorithms like RLHF and DPO operate as single optimization procedures.
@@
 (2024a) proposed that LLMs can refine their capabilities through self-play by interacting with instances of themselves.
-More specifically, the LLM generates its own training data from prior iterations and subsequently learns a new policy that outperforms the old one (Wu et al., 2024a).
-This iterative fine-tuning framework has been shown to be equivalent to finding the Nash equilibrium of a two-player game (Munos et al., 2023; Rosset et al., 2024; Calandriello et al., 2024), gaining significant recognition due to its solid theoretical foundation and simplicity (Gao et al., 2024a; Ye et al., 2024).
+More specifically, the LLM generates its own training data from prior iterations and subsequently learns a new policy that outperforms the old one (Wu et al., 2025).
+This iterative fine-tuning framework has been shown to be equivalent to finding the Nash equilibrium of a two-player game, gaining significant recognition due to its solid theoretical foundation and simplicity (Calandriello et al., 2024).
 In recommender systems, we leverage the self-play mechanism to adaptively reduce biased items in a simple and effective manner.
@@
 In this section, we provide a brief overview of the technologies for aligning LLMs with the recommendation task. We then introduce the idea of evaluating the biases and unfairness in LRSs.
-Figure 2. Illustration of the pipelines and principles of SFT, DPO, and our proposed SPRec in LLM-based recommendations.
 3.1. Supervised Fine-tuning (SFT)
-To enable an open-sourced LLM to effectively learn recommendation tasks, one of the most practical approaches is fine-tuning all or a subset of its parameters using a set of demonstration data derived from offline recommendation logs.
-The fine-tuning process aligns the model’s behavior with the recommendation objective by optimizing its parameters on the given dataset.
-The objective function for this fine-tuning process is defined as maximizing the log-likelihood over the offline training dataset $\mathcal{D}$:
+To enable an open-source LLM to learn recommendation tasks effectively, a practical approach is fine-tuning all or part of its parameters using demonstration data from offline recommendation logs.
+The objective is to align the model’s behavior with the recommendation task by maximizing the log-likelihood over the training dataset $\mathcal{D}$:
 (1) $\pi_{SFT}=\arg\max_{\pi_{\theta}}\mathbb{E}_{(x_{i},y_{i})\sim\mathcal{D}}\log\pi_{\theta}(y_{i}|x_{i}),$
-where $(x_{i},y_{i})$ denotes the input-output pairs sampled from the training dataset $\mathcal{D}$.
-Here, $x_{i}$ represents the context of a recommendation, typically comprising user features and a sequence of historically interacted items, while $y_{i}$ denotes the target item to be recommended.
-Let $p_{\mathcal{D}}(y|x)$ be the probability of $y$ given $x$ in the dataset $\mathcal{D}$, i.e., the popularity.
-SFT’s formulation is equal to learning to align the model’s predictions with it by minimizing the forward KL-divergence $\mathbb{D}_{KL}(p_{\mathcal{D}}(y|x),\pi_{\theta}(y|x))$:
+where $(x_{i},y_{i})$ are input-output pairs from $\mathcal{D}$, with $x_{i}$ representing user context and interaction history, and $y_{i}$ the target item.
+Defining $p_{\mathcal{D}}(y|x)$ as the empirical probability (i.e., item popularity), SFT aligns model predictions by minimizing the forward KL-divergence:
 $\pi_{SFT}$ $=\arg\min_{\pi_{\theta}}\mathbb{D}_{KL}(p_{\mathcal{D}}(y|x),\pi_{\theta}(y|x))$
-(2) $=\arg\min_{\pi_{\theta}}\mathbb{E}_{(x_{i},y_{i})\sim D}-\log\pi_{\theta}(y_{i}|x_{i})+H(p_{\mathcal{D}}),$
-where $H(p_{\mathcal{D}})$ denotes the entropy of $p_{\mathcal{D}}$, which is constant with respect to the optimization of the LLM parameters $\theta$.
-However, relying exclusively on subsequent targets as positive samples may confuse the model about the true objectives.
-In recommendation tasks, negative samples play a critical role in highlighting undesirable outcomes (Shi et al., 2023).
-Thus, it is natural to consider incorporating both positive and negative samples from the offline data to better align the model with user preferences.
-3.2. RLHF and DPO
-To ensure that outputs align with intricate user preferences (such as helpfulness and harmlessness (Bai et al., 2022)), researchers have introduced Reinforcement Learning with Human Feedback (RLHF) to train LLMs using reward signals (Ouyang et al., 2022).
-Typically, the RLHF stage follows the SFT stage, further refining language models to better adhere to human instructions.
-RLHF involves learning a reward model $r_{\phi}(x,y)$, which captures human preferences.
-The optimization objective during this stage is defined as:
-(3) $\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(y|x)}[r_{\phi}(x,y)]-\beta\mathbb{D}_{KL}[\pi_{\theta}(y|x),\pi_{\mathrm{ref}}(y|x)],$
-where $\beta$ is a regularization term that penalizes deviations of the learned model $\pi_{\theta}$ from the reference policy $\pi_{\mathrm{ref}}$.
-While RLHF has demonstrated remarkable success in aligning LLMs with human preferences, it is computationally intensive, complex, and heavily reliant on the reward model for effectiveness.
-To simplify reward modeling and mitigate the instability associated with RLHF training, researchers have proposed Direct Preference Optimization (DPO) (Rafailov et al., 2024), which utilizes a closed-form solution of RLHF’s optimal policy.
-The resulting objective function for DPO is:
+(2) $=\arg\min_{\pi_{\theta}}\mathbb{E}_{(x_{i},y_{i})\sim\mathcal{D}}-\log\pi_{\theta}(y_{i}|x_{i})+H(p_{\mathcal{D}}),$
+where $H(p_{\mathcal{D}})$ is the constant entropy of $p_{\mathcal{D}}$.
+3.2. Direct Preference Optimization (DPO)
+To ensure that model outputs align with intricate user preferences, researchers have proposed Direct Preference Optimization (DPO) (Rafailov et al., 2024), which optimizes the following objective function:
 (3) $\min_{\pi_{\theta}}-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\log\sigma\Bigg{[}\!\beta\log\!\left(\!\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\mathrm{ref}}(y_{w}|x)}\!\right)-\beta\log\!\left(\!\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\mathrm{ref}}(y_{l}|x)}\!\right)\!\Bigg{]},$
-where $(x,y_{w},y_{l})$ represents a prompt $x$ alongside a chosen (preferred) answer $y_{w}$ and a rejected answer $y_{l}$.
-In the context of recommendation tasks, $x$ corresponds to the recommendation context, as in the SFT stage, while $y_{w}$ and $y_{l}$ denote a positive and a negative sample, respectively.
-DPO offers an efficient and stable alternative to RLHF, aligning models with user preferences while reducing reliance on complex reward modeling.
-Its ability to naturally leverage positive and negative samples makes it particularly well-suited for recommender systems (Chen et al., 2024b; Bai et al., 2024; Liao et al., 2024a).
-3.3. Evaluating Biases and Unfairness through Distribution Alignment Perspective
+where $(x,y_{w},y_{l})$ denotes a prompt $x$ with a chosen (preferred) answer $y_{w}$ and a rejected answer $y_{l}$.
+The parameter $\beta$ acts as a regularization factor, controlling the extent to which the learned policy $\pi_{\theta}$ deviates from the reference policy $\pi_{\mathrm{ref}}$.
+In the context of recommendation tasks, $x$ represents the user context, typically comprising user features and historical interaction sequences, while $y_{w}$ and $y_{l}$ correspond to positive and negative samples, respectively.
+The goal is to encourage the model to assign higher probabilities to preferred items ($y_{w}$) over less desirable ones ($y_{l}$), effectively capturing user preferences.
+DPO offers an efficient and stable solution for preference alignment, eliminating the need for complex reward models often required in reinforcement learning-based approaches.
+Its natural ability to incorporate both positive and negative samples makes it particularly well-suited for recommender systems, where learning from contrasting user interactions is crucial (Chen et al., 2024b; Bai et al., 2024; Liao et al., 2024).
+3.3. Evaluating Bias via Distribution Alignment
 When aligning user preferences, LLMs may inadvertently learn biased or unfair outcomes.
@@
 Specifically, let $R$ denote the ground-truth user preference (e.g., an item list), following the distribution $P(R)$, and let $\hat{R}$ represent the model-predicted preferences, drawn from $P(\hat{R})$.
-Bias or unfairness is then quantified by the mismatch between these two distributions:
-(5) $P(R)\neq P(\hat{R}).$
-To operationalize this framework, we adopt the method outlined by Jiang et al.
-(2024), where $P(R)$, the target user history preference distribution, is approximated using the category distribution derived from offline training data.
-Furthermore, we also employ their proposed metrics, MGU@$K$ and DGU@$K$, to systematically measure the degree of mismatch in our experiments.
-4. Method
-In this section, we first point out the severe popularity bias due to the nature of DPO, then we analyze how we can address this by utilizing the self-play philosophy.
-Then we detail the proposed SPRec architecture.
-4.1. Problem of DPO: Amplify Popular Bias
-As we mentioned DPO is effective and efficient in aligning LLMs with the user preferences, it has a problem of suppressing the negative samples too much, which will result in severe popularity bias in the recommendation task.
+Bias or unfairness is then quantified by the mismatch between these two distributions: $P(R)\neq P(\hat{R})$.
+To apply this framework, we follow Jiang et al.
+(2024), approximating $P(R)$ using the category distribution from offline training data.
+We further employ their MGU metric to systematically measure the degree of mismatch in our experiments.
+4. Problem of DPO: Amplify Popularity Bias
+We present an empirical analysis to demonstrate how DPO exacerbates popularity bias in LRSs, followed by a theoretical examination of the underlying mechanisms driving this phenomenon.
+Figure 2. Distribution of cold-start recommendation results. Group 0: least popular, group 4: most popular.
+Figure 3.
+Illustration of SFT, DPO, and SPRec in LLM-based recommendations.
+(a) SFT generates mass-covering results but retains inherent biases.
+(b) DPO amplifies these biases by over-representing certain items.
+(c) SPRec mitigates bias through self-play, leveraging model outputs as negative samples to achieve balanced recommendations.
+4.1. Empirical Analysis
+To systematically evaluate how DPO amplifies recommendation bias, we design a cold-start recommendation task using two widely-used benchmark datasets: MovieLens and Goodreads [^1] Dataset details are provided in Section 6.1.1.
+In this task, the LLM generates recommendations without access to user interaction history.
+We randomly sample 100 items and partition them into five groups based on interaction probabilities, ensuring balanced popularity levels.
+Positive samples are drawn accordingly, while negative samples for DPO training are randomly selected.
+The training and validation sets contain 4096 and 512 samples, respectively.
+After training, we analyze the distribution of recommendations across the five groups.
+Fig.
+2 shows the proportion of recommendations allocated to each group before and after training, where group 0 represents the least popular items and group 4 the most popular.
+The results reveal three key insights: (1) SFT introduces a slight bias toward group 4; (2) DPO significantly amplifies this bias, causing recommendations to concentrate almost entirely on the most popular items; and (3) our proposed method, SPRec (introduced in detail later), effectively mitigates the bias amplification induced by DPO.
+4.2. Theoretical Analysis
 In recommendation tasks, the input and positive samples $(x,y_{w})$ are derived from logged interactions in offline data, while negative samples $y_{l}$ are drawn from non-interacted items.
-Given an input $x$, the probability $p_{\mathcal{D}}(y|x)$ represents the popularity of item $y$ in the dataset, and $q_{\mathcal{D}}(y|x)$ denotes the probability of the same item $y$ being selected as a negative sample.
+Given an input $x$, the probability $p_{\mathcal{D}}(y|x)$ represents the conditional popularity of item $y$ in the dataset, and $q_{\mathcal{D}}(y|x)$ denotes the probability of the same item $y$ being selected as a negative sample.
 Using these definitions, the DPO loss in Eq.
@@
 This result highlights that the optimal policy is proportional to the reference policy $\pi_{\mathrm{ref}}(y|x)$, adjusted by the relative likelihood ratio $\left(\frac{p_{\mathcal{D}}(y|x)}{q_{\mathcal{D}}(y|x)}\right)^{1/\beta}$.
-In most recommendation setting, the probability of an item $y$ being a negative sample satisfies a uniform distribution (Bai et al., 2024; Chen et al., 2024b), i.e., $q_{\mathcal{D}}(y|x)=\mathcal{U}=\frac{1}{|\mathcal{I}|}$, where $\mathcal{I}$ is the all item candidates.
-Furthermore, in most preference alignment scenarios using the DPO-based method, the parameter $\beta$ is constrained to the range $10<\beta<$ 1.
-Under these conditions, the DPO loss naturally encourages the policy to prioritize popular items with higher probabilities $p_{\mathcal{D}}(y|x)$, thereby introducing a notable popularity bias.
-In the extreme case where $\beta\to 0$, the optimal policy under the DPO loss degenerates to recommending only the most popular items with the highest probabilities $p_{\mathcal{D}}(y|x)$, effectively excluding less popular but potentially valuable items from consideration.
+In most recommendation settings, negative samples are uniformly distributed (Bai et al., 2024; Chen et al., 2024b), i.e., $q_{\mathcal{D}}(y|x)=\mathcal{U}=\frac{1}{|\mathcal{I}|}$, where $\mathcal{I}$ is the set of all candidate items.
+Additionally, in typical DPO-based preference alignment scenarios, $\beta$ is constrained to $10<\beta<$ 1.
+Under these conditions, the DPO loss inherently biases the model toward popular items with higher $p_{\mathcal{D}}(y|x)$, exacerbating popularity bias.
+In the extreme case where $\beta\to 0$, the optimal policy collapses to recommending only the most popular items, effectively disregarding less frequent but potentially valuable recommendations.
 Remark: This result is a byproduct of DPO’s loss function.
@@
 Forward KL-divergence is known for its mass-covering property, which encourages learning an average behavior and is less sensitive to subtle differences in the preference distribution (as illustrated in Fig.
-2(a)) (Sun and van der Schaar, 2024).
-In contrast, the reverse KL-divergence used in DPO promotes mode-seeking behavior (Wang et al., 2023a; Omura et al., 2024), guiding the model to focus on the “peaks” of the distribution (Fig.
-2(b)).
+3(a)) (Sun and van der Schaar, 2024).
+In contrast, the reverse KL-divergence used in DPO promotes mode-seeking behavior (Wang et al., 2024a; Omura et al., 2024), guiding the model to focus on the “peaks” of the distribution (Fig.
+3(b)).
 The issue of DPO has also been highlighted by other researchers (Pal et al., 2024).
@@
 In the context of recommendation, this behavior can be detrimental, as it exacerbates the filter bubble issue and undermines user interests by limiting exposure to diverse items (Gao et al., 2023a).
+5. Method
+We present how to address the popularity bias in DPO by utilizing the self-play philosophy. Then we detail the proposed SPRec architecture.
 5.1. Solution: Suppress Biases through Self-Play
-Since the DPO loss inherently causes the policy $\pi_{\theta}$ to learn sharp ”peaks”, leading to bias, an intuitive solution is to directly suppress these learned peaks.
+Since the DPO loss inherently causes the policy $\pi_{\theta}$ to learn sharp “peaks”, leading to bias, an intuitive solution is to directly suppress these learned peaks.
 To address this, we utilize a self-play framework, dubbed SPDPO, which iteratively alternates between policy learning and bias suppression.
@@
 The workflow is illustrated in Fig.
-2(c), in which three key steps are conducted sequentially in each iteration:
+3(c), in which three key steps are conducted sequentially in each iteration:
   1. (1)
@@
     DPO Step: Align the model $\pi_{\theta_{t}}$ by perform DPO step using pairwise dataset $\{(x^{i},y_{w}^{i},y_{l}^{i})\}$, and obtain $\pi_{\theta_{t+1}}$.
-The iteration will repeat $T$ times in an epoch.
-The self-play philosophy can utilized in any LLM tuning-based LRSs.
-Next, we will explain the necessity of the SFT step, which is not considered in original self-play fine-tuning work (Chen et al., 2024a).
+This process repeats for $T$ iterations per epoch.
+The self-play mechanism is adaptable to any LLM-based recommender system.
+To ensure comparability with existing DPO-based recommenders (Chen et al., 2024b; Bai et al., 2024; Liao et al., 2024), we can extend from a single to multiple negative samples, with results analyzed in experiments.
 Remark: Although the loss function in Eq.
@@
 Similarly, Pang et al.
-(2024) demonstrate the effectiveness of an iterative preference optimization algorithm useing a modified DPO loss with an additional negative log-likelihood (NLL) term, which mirrors the SFT loss defined in Eq.
+(2024) demonstrate the effectiveness of an iterative preference optimization algorithm using a modified DPO loss with an additional negative log-likelihood (NLL) term, which mirrors the SFT loss defined in Eq.
 (1).
-Table 1. Statistics of Datasets.
-Datasets #Items #Interactions #Sequences
-MovieLens 10,682 10,000,054 9,301,274
-Goodreads 4,058 160,398 6,031
-CDs_and_Vinyl 13,078 185,855 21,347
-Steam 32,094 178,961 29,876
 6. Experiments
@@
 Table 1.
-Overall performance comparison of SPRec (green) with SFT-based (brown) and DPO-based (blue) methods.
-Best results are in bold, sub-optimal results are underlined.
-Metrics with $\uparrow$ indicate higher is better, while $\downarrow$ indicates lower is better.
-Dataset Model DivRatio@1$\uparrow$ DivRatio@5$\uparrow$ ORRatio@1$\downarrow$ ORRatio@5$\downarrow$ MGU@1$\downarrow$ MGU@5$\downarrow$ HR@5$\uparrow$ NDCG@5$\uparrow$
-SASRec 0.0133 0.0057 0.6588 0.4778 0.0463 0.0439 0.0240 0.0155
+Overall performance comparison of SPRec (green), SFT-based (brown), and DPO-based (blue) methods.
+Best results are bold, sub-optimal ones underlined.
+$\uparrow$ indicates higher is better, while $\downarrow$ indicates lower is better.
 […]
-RW 0.1918 0.1541 0.2551 0.0990 0.0577 0.0405 0.0327 0.0276
-$\text{D}^{3}$ 0.1246 0.1000 0.3238 0.1440 0.0664 0.0502 0.0266 0.0197
+SASRec 0.0031 1.0000 0.1209 0.0225 0.0136
+[…]
+RW 0.1918 0.2551 0.0577 0.0327 0.0276
+$\text{D}^{3}$ 0.1246 0.3238 0.0664 0.0266 0.0197
 […]
@@
 […]
-SASRec 0.3236 0.1889 0.1391 0.1149 0.0070 0.0063 0.1149 0.0837
+SASRec 0.0030 1.0000 0.0458 0.0202 0.0139
 […]
-RW 0.1050 0.0820 0.3730 0.1338 0.0238 0.0270 0.0380 0.0281
+RW 0.1050 0.3730 0.0238 0.0380 0.0281
 […]
@@
 […]
-SASRec 0.0021 0.0035 1.0000 0.5529 0.1446 0.0632 0.0230 0.0161
+SASRec 0.0021 1.0000 0.1205 0.0226 0.0170
 […]
-RW 0.2821 0.2646 0.2648 0.0971 0.0432 0.0305 0.0112 0.0103
-$\text{D}^{3}$ 0.3135 0.2968 0.2672 0.0748 0.0320 0.0252 0.0103 0.0088
+RW 0.2821 0.2648 0.0432 0.0112 0.0103
+$\text{D}^{3}$ 0.3135 0.2672 0.0320 0.0103 0.0088
+[…]
+SDPO 0.3218 0.2118 0.0380 0.0110 0.0106
 […]
 […]
+SASRec 0.0010 1.0000 0.1094 0.0660 0.0379
 […]
 […]
-SASRec 0.3640 0.1794 0.1610 0.1102 0.0241 0.0209 0.1940 0.1407
+$\text{D}^{3}$ 0.1580 0.4560 0.0511 0.0730 0.0718
 […]
-[…]
-$\text{D}^{3}$ 0.1580 0.1414 0.4560 0.1896 0.0511 0.0320 0.0730 0.0718
-[…]
-SDPO 0.2140 0.1978 0.3270 0.0756 0.0448 0.0320 0.0810 0.0804
-RosePO 0.2060 0.1868 0.3670 0.1080 0.0562 0.0349 0.0760 0.0750
+SDPO 0.2080 0.3510 0.0475 0.0820 0.0810
+RosePO 0.2310 0.3160 0.0499 0.0820 0.0805
 […]
@@
 6.1.1. Datasets
-We conducted extensive experiments on four real-world datasets: MovieLens [^1] https://grouplens.org/datasets/movielens/ (Harper and Konstan, 2015), Steam [^2] https://cseweb.ucsd.edu/j̃mcauley/datasets.html#amazon_reviews, Goodreads [^3] https://mengtingwan.github.io/data/goodreads, and the CDs and Vinyl category of the Amazon Review Dataset [^4] http://jmcauley.ucsd.edu/data/amazon/index_2014.html.
+We conducted extensive experiments on four real-world datasets: MovieLens [^2] https://grouplens.org/datasets/movielens/, Steam [^3] https://cseweb.ucsd.edu/j̃mcauley/datasets.html#amazon_reviews, Goodreads [^4] https://mengtingwan.github.io/data/goodreads, and the CDs and Vinyl category of the Amazon Review Dataset [^5] http://jmcauley.ucsd.edu/data/amazon/index_2014.html.
+Additional details about the datasets are provided in the Appendix B.
 Following the data processing approach in (Chen et al., 2024b; Bao et al., 2024), interaction sequences with fewer than 10 entries were excluded.
 The datasets were then split chronologically into training, validation, and test sets in an 8:1:1 ratio, ensuring mutual exclusivity and preventing data leakage.
-The statistics of the datasets are summarized in Table 1.
+To ensure comparability across different LLM-based methods, we further sampled 4,096 interactions from each dataset’s training set as the training samples for all methods, 512 interactions from the validation set, and 1,000 interactions from the test set.
 To process category information, we extracted category metadata from each dataset and identified the most $10$ popular categories within the training sets.
 To ensure category independence, we removed categories with clear hierarchical relationships, such as “FPS” and “Shooting” in the Steam dataset, and “Rock” and “Classical Rock” in the CDs and Vinyl dataset.
-Table 3. Ablation Experiment Results. SPRec-RN represents the training process with negative samples randomly sampled, while SPRec-noDPO and SPRec-noSFT denote SPRec without DPO and SFT, respectively.
-Dataset Model DivRatio@1$\uparrow$ DivRatio@5$\uparrow$ ORRatio@1$\downarrow$ ORRatio@5$\downarrow$ MGU@1$\downarrow$ MGU@5$\downarrow$ HR@5$\uparrow$ NDCG@5$\uparrow$
-SPRec-noSFT 0.3020 0.2486 0.0837 0.0224 0.0198 0.0188 0.0184 0.0149
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-SPRec-noDPO 0.2220 0.2006 0.3910 0.1884 0.0620 0.0294 0.0790 0.0776
-[…]
-[…]
 Figure 4. Comparison of models across genres on Group Unfairness (GU) in top-$1$ recommendation.
@@
 To leverage the strengths of LLMs in generative recommendation tasks, we prompt the LLM to generate a predicted item based on the input history sequence.
-Then, following the procedures in BIGRec (Bao et al., 2023a), we calculate scores and rankings for the entire item space and ground our predicted item to an exact item in the dataset.
+Then, following the procedures in BIGRec (Bao et al., 2025), we calculate scores and rankings for the entire item space and ground our predicted item to an exact item in the dataset.
 6.1.3. Metrics.
-We evaluate the model using four key metrics on a randomly sampled set of 1,000 test cases per iteration.
-Accuracy is measured using Normalized Discounted Cumulative Gain (NDCG@$k$) and Hit Ratio (HR@$k$) and calculated per result before averaging.
-Diversity is assessed with DivRatio@$k$, representing the proportion of unique items in the top-$k$ recommendations.
-Over-recommendation is quantified by ORRatio@$k$, indicating the proportion of the top-$k$ results consisting of the three most frequently recommended items.
-Fairness is evaluated using MGU@$k$ (Jiang et al., 2024), which captures category-level discrepancies between the model’s top-$k$ recommendations and the user’s historical sequences.
+We evaluate the model on 1,000 randomly sampled test cases per iteration using four key metrics.
+Accuracy is measured by NDCG@5 and HR@5, averaged across results.
+Diversity is assessed via DivRatio, representing the proportion of unique recommendations.
+Over-recommendation is quantified by ORRatio, indicating the proportion of results dominated by the three most frequently recommended items.
+Fairness is evaluated using MGU (Jiang et al., 2024), capturing category-level discrepancies between recommendations and user history.
 6.1.4. Baseline
@@
 For LLM-based models, we consider several baselines.
-(1) For SFT-based methods, BIGRec (Bao et al., 2023a) serves as an instruction-tuning LLM framework for sequential recommendations and forms the foundation for SPRec.
+(1) For SFT-based methods, BIGRec (Bao et al., 2025) serves as an instruction-tuning LLM framework for sequential recommendations and forms the foundation for SPRec.
 Re-weighting (RW) (Jiang et al., 2024) improves fairness in BIGRec by balancing recommendations across categories through dataset-based training weights.
@@
 (2) For DPO-based models, DMPO (Bai et al., 2024) introduces DPO into LRSs by sampling multiple negative items as rejected responses, while Softmax-DPO (SDPO) (Chen et al., 2024b) follows a similar approach but incorporates a softmax loss over multiple negative samples.
-Finally, RosePO (Liao et al., 2024a) is a preference optimization framework that combines negative sampling strategies and personalized uncertainty to achieve fairness, unbiasedness, and robustness.
-The implementation details are listed in Appendix B.
+Finally, RosePO (Liao et al., 2024) is a preference optimization framework that combines negative sampling strategies and personalized uncertainty to achieve fairness, unbiasedness, and robustness.
+The implementation details are listed in Appendix C.
 6.2. Overall Performance Comparison (RQ1)
-The experimental results are shown in Table 5, leading to the following observations.
+The experimental results are presented in Table 6, leading to the following observations.
+The non-LLM baseline, SASRec, performs poorly with the given training size, which is expected as SASRec requires large datasets to achieve effective fitting.
+In this study, we primarily focus on LLM-based methods, and SASRec’s results are included only for reference and as the assistant model for $\text{D}^{3}$ during the decoding stage.
 6.2.1. Limitations of SFT-based Methods
 Fine-tuning LLMs with instruction-based methods results in recommendations heavily favoring popular items, leading to a lack of diversity.
-For example, in the Goodreads dataset, the DivRatio@1 of BIGRec is only 0.142, meaning the model provides just 14 distinct recommendations per 100 tasks.
-Similarly, in the Steam dataset, BIGRec’s ORRatio@1 reaches 0.391, with over 39% of recommendations concentrated on the 3 most popular items.
+For example, in the Goodreads dataset, the DivRatio of BIGRec is only 0.142, meaning the model provides just 14 distinct recommendations per 100 tasks.
+Similarly, in the Steam dataset, BIGRec’s ORRatio reaches 0.391, with over 39% of recommendations concentrated on the 3 most popular items.
 These findings highlight that relying solely on SFT introduces severe biases, significantly overexposing certain popular items.
@@
 For DPO methods using random sampling, such as SDPO and DMPO, while multiple negative samples improve recommendation accuracy, they perform poorly on diversity and fairness metrics.
-In the Goodreads and MovieLens datasets, SDPO and DMPO exhibit significantly higher ORRatio values than BIGRec, supporting the conclusion in Section 4.1 that DPO amplifies popularity bias.
-On the CDs and Steam datasets, although ORRatio improves, diversity metrics remain largely unchanged, indicating that the model favors moderately popular items but fails to effectively recommend new ones.
-In summary, existing DPO-based methods not only fail to address fairness issues in LRS but may exacerbate popularity bias.
+On the Goodreads and MovieLens datasets, SDPO and DMPO have minimal impact on DivRatio and ORRatio and may even degrade model performance.
+On the CDs and Steam datasets, although ORRatio decreases, diversity metrics remain largely unchanged, suggesting that the model favors moderately popular items but fails to effectively recommend new ones.
+In contrast, RosePO performs well on the CD dataset due to its negative sampling strategy based on semantic information.
+However, this approach heavily relies on the semantic characteristics of the dataset’s structure, resulting in relatively poor performance on other datasets and limiting its generalizability for debiasing.
+In summary, existing DPO-based methods fail to address fairness issues in LRS.
 6.2.3. Superiority of SPRec
@@
 Additionally, SPRec outperforms BIGRec on most fairness metrics, reducing the discrepancies between the model’s recommendations and users’ historical sequences, thereby providing fairer recommendations.
-SPRec also surpasses most baseline models on DivRatio and ORRatio across the majority of datasets, showcasing its superior ability to balance recommendation distributions.
-For fairness, SPRec achieves the best MGU@1 on all datasets except Steam.
+SPRec also surpasses all baseline models on DivRatio and ORRatio, showcasing its superior ability to balance recommendation distributions.
+For fairness, SPRec achieved the highest MGU scores on the MovieLens and Goodreads datasets, and the second-highest on the Steam and CD datasets.
 Moreover, as shown in Fig.
-3, SPRec alleviates category-level unfairness on the MovieLens dataset, achieving the best results in 7 out of 8 categories, further underscoring its effectiveness in improving fairness.
+4, SPRec alleviates category-level unfairness on the MovieLens dataset, achieving the best results in 7 out of 8 categories, further underscoring its effectiveness in improving fairness.
 While RosePO performs well on the CDs and Vinyl dataset, leveraging semantic-based negative sampling to address fairness in music recommendations, and Re-weighting shows strong performance on the Steam dataset by employing category-based re-weighting for gaming recommendations, these methods are tailored to specific datasets and lack generalizability.
 In contrast, SPRec’s self-play framework provides a universal solution, overcoming dataset-specific challenges and delivering fairer recommendations across diverse scenarios.
+Table 2. Ablation results. “with RN” for random negative samples, “w/o” for without specific components.
+[…]
+w/o SFT 0.3020 0.0837 0.0198 0.0184 0.0149
+[…]
+with RN 0.2194 0.2224 0.0544 0.0286 0.0230
+[…]
+[…]
+[…]
+with RN 0.1380 0.3380 0.0188 0.0420 0.0310
+[…]
+w/o SFT 0.3381 0.2373 0.0216 0.0132 0.0126
+[…]
+[…]
+[…]
+[…]
+w/o DPO 0.2220 0.3910 0.0620 0.0790 0.0776
+[…]
+Steam SPRec 0.2930 0.2560 0.0367 0.0910 0.0893
 Figure 5. Performance on the MovieLens dataset across different ablation experiments.
@@
 6.3.1. SPRec without SFT
-As shown in Table 3, SPRec-noSFT achieves the lowest recommendation accuracy across all datasets except Steam.
+As shown in Table 2, SPRec w/o SFT achieves the lowest recommendation accuracy across all datasets except Steam.
 This indicates that, during the self-play process, the model’s excessive focus on fairness compromises its accuracy.
 In the MovieLens dataset (Fig.
-4), the absence of the SFT stage leads to a steady decline in recommendation accuracy (NDCG@1) throughout training.
+5), the absence of the SFT stage leads to a steady decline in recommendation accuracy (NDCG) throughout training.
 These findings highlight the critical role of SFT in maintaining SPRec’s recommendation performance.
 6.3.2. SPRec without DPO
-Removing DPO reduces SPRec to further SFT training on new data, ensuring that performance gains are not due to incorporating additional data.
-As shown in Table 3, even with new data, additional SFT fails to improve diversity or fairness metrics, and the recommendations remain biased.
+Removing DPO reduces SPRec to further SFT training, ensuring that performance gains are not due to incorporating additional data.
+As shown in Table 2, additional SFT fails to improve diversity or fairness metrics, and the recommendations remain biased.
 Furthermore, Fig.
@@
 6.3.3. Randomly sampling negative items
+As observed in Table 2, when the negative sampling strategy is replaced with random sampling, SPRec-RN fails to achieve further improvements in DivRatio and ORRatio metrics on the MovieLens and Goodreads datasets.
+Additionally, SPRec-RN’s fairness metrics perform worse compared to SPRec.
+Although SPRec-RN shows a significant improvement in DivRatio on the CDs and Vinyl dataset, its ORRatio still performs poorly.
+This suggests that random sampling of negative samples during training is ineffective at suppressing popular items, and the recommendation results continue to exhibit a significant long-tail effect.
+This ablation experiment demonstrates that our Self-play negative sampling strategy effectively balances the distribution of the model’s output, leading to debiasing in recommendations.
 Replacing the negative sampling strategy with random sampling (SPRec-RN) fails to improve DivRatio and ORRatio metrics on the MovieLens and Goodreads datasets (Table 2).
@@
 To ensure result stability, the model’s performance is averaged over the last three training iterations.
+We report the results on the Movielens dataset.
 6.4.1. Effect of Random Sampling Ratio
@@
 As shown in Fig.
-5, increasing the proportion of random negative samples leads to a steady decline in diversity and accuracy.
+6, increasing the proportion of random negative samples leads to a steady decline in diversity and accuracy.
 Fairness also deteriorates, with recommendations becoming more skewed toward a small subset of popular items.
@@
 As shown in Fig.
-6, increasing the number of negative samples results in stable recommendation accuracy but significantly improves diversity and fairness, reducing the focus on popular items.
+7, increasing the number of negative samples results in stable recommendation accuracy but significantly improves diversity and fairness, reducing the focus on popular items.
 This demonstrates the versatility of the self-play negative sampling strategy, which can be effectively combined with multi-negative sampling approaches to further debias LRS.
-6. Conclusion
-This paper theoretically and experimentally analyzes the limitations of existing SFT- and DPO-based methods in LLM-based recommender systems.
-These approaches often promote a narrow set of items, intensifying the filter bubble effect, harming user interests, and raising fairness concerns that may undermine creators’ motivation.
-To address these challenges, we propose SPRec, a novel self-play recommendation framework that mitigates over-recommendation and enhances fairness without requiring additional data or manual intervention.
-SPRec leverages the model’s output from the previous round as negative samples, effectively re-weighting the DPO loss function with the model’s logits.
-This adaptive mechanism suppresses biased items and improves recommendation quality.
-We conducted experiments on four real-world datasets, validating the effectiveness of our method in improving recommendation accuracy, diversity, and fairness.
-Comprehensive ablation studies revealed that fairness improvements largely stem from the suppression of self-play negative samples during DPO.
-In the future, we plan to explore more advanced preference alignment techniques for LRSs, aiming for smarter personalized decision-making and more helpful and harmless recommendation systems.
+7. Conclusion & Discussion
+Our work establishes a critical bridge between preference alignment techniques and fairness-aware recommendation in the era of LLMs.
+Through both theoretical analysis and empirical validation, we demonstrate that conventional DPO-based tuning fundamentally conflicts with the principles of equitable recommendation, creating self-reinforcing popularity biases that traditional debiasing approaches fail to address.
+The proposed SPRec framework represents a paradigm shift in recommendation alignment - rather than treating bias mitigation as a post-hoc correction, we redesign the core learning mechanism to enable autonomous bias suppression through self-regulated competition between model generations.
+This approach not only achieves state-of-the-art performance across accuracy and fairness metrics but more importantly, provides a blueprint for developing self-correcting AI systems that maintain alignment with both user preferences and ethical constraints.
+Despite its effectiveness, our work primarily addresses bias in DPO-based tuning, while overlooking the popularity bias already present in SFT due to its cross-entropy loss.
+Future research should focus on mitigating bias at the SFT stage to ensure fairness from the start of fine-tuning.
+Additionally, optimizing preferences in recommendation is a long-term challenge, requiring alignment across sequential recommendations rather than individual predictions.
+However, LLMs generate outputs token by token, making it difficult to optimize preferences from token-level to item-level and ultimately list-level recommendations.
+Tackling this issue will require new datasets, benchmarks, and models capable of long-term alignment.
+A promising direction is reinforcement learning with process-level rewards, shifting optimization from short-term token likelihood to long-horizon user engagement.
+Acknowledgements
+This work is supported by the National Key Research and Development Program of China (2021ZD0111802), the National Natural Science Foundation of China (62402470, 62272437, U24B20180, 62121002), the Fundamental Research Funds for the Central Universities of China (WK2100000053, PA2024GDSK0107), Anhui Provincial Natural Science Foundation (2408085QF189), and the Postdoctoral Fellowship Program of CPSF (GZC20241643).
+This research is supported by the advanced computing resources provided by the Supercomputing Center of the USTC.
 Appendix A Mathematical Derivations
 Proof of Theorem 1.
-The DPO loss is derived from the RLHF objective as follows:
+The DPO loss is derived from the objective of Reinforcement Learning with Human Feedback (RLHF):
 (7) $\max_{\theta}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(\cdot|x)}[r(x,y)]-\beta\mathrm{D_{KL}}(\pi_{\theta}\|\pi_{\mathrm{ref}}),$
@@
 In the original paper of DPO (Rafailov et al., 2024), the authors proved that the optimal policy $\pi_{\theta}^{*}$ for DPO loss in Eq.
-(LABEL:eqn:dpo_def) and the solution to the optimization problem in Eq.
+Eq.
+(4) and the solution to the optimization problem in Eq.
 (7) are the same.
@@
 $\max_{\theta}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(\cdot|x)}[\log p_{\mathcal{D}}(y|x)-\log q_{\mathcal{D}}(y|x)]-\beta\mathrm{D_{KL}}(\pi_{\theta}\|\pi_{\mathrm{ref}}),$
-which has a well-known closed form solution (Rafailov et al., 2024): π_θ^*(y—x) ∝π_ref(y—x)⋅(pD(y—x)qD(y—x))^1/β. ∎
-Appendix B Implementation Details
-For all traditional models, we set the sizes of the training and validation sets to $65,536$ and $8,192$, respectively.
+which has a well-known closed form solution (Rafailov et al., 2024): π_θ^*(y—x) ∝π_ref(y—x)⋅(pD(y—x)qD(y—x))^1/β.
+Appendix B Dataset Statistics
+Our datasets span diverse domains, including movies, books, music, and games, offering varied sizes and user interaction patterns to provide a comprehensive basis for evaluating LRSs.
+Note that while we report the full dataset statistics, only a subset of interaction sequences is sampled for LLM fine-tuning, as detailed in Section 6.1.1.
+Table 3. Statistics of Datasets.
+Datasets #Items #Interactions #Sequences
+MovieLens 10,682 10,000,054 9,301,274
+Goodreads 4,058 160,398 6,031
+CDs_and_Vinyl 13,078 185,855 21,347
+Steam 32,094 178,961 29,876
+Appendix C Implementation Details
+For LLM-based methods, we adopted Llama-3.2-1B-Instruct as the backbone LLM.
+Considering the ability of LLMs to quickly adapt to downstream tasks with limited data, we followed BIGRec (Bao et al., 2025) and used relatively smaller datasets.
+To ensure fairness in comparison, all baseline methods and SPRec utilize the same dataset as used in the SFT training phase.
+For SPRec, the total number of iterations was set to $5$, with each SFT and DPO phase trained for 1 epoch.
+To ensure that the training data used in each iteration is not identical, we further randomly sample half of the training data (i.e., 2048 interactions) for training in each iteration.
+All experiments were carried out on four RTX 3090 GPUs, each with 24GB of VRAM.
+For the traditional model SASRec, we use the same training and validation datasets as other LLM-based methods, with dataset sizes of 4,096 and 512, respectively.
 The embedding size was fixed at $64$, and the dropout ratio was set to 0.1.
 Negative samples were randomly sampled in training, with Adam as the optimizer and a learning rate of 4e-3.
-For LLM-based methods, we adopted Llama-3.2-1B-Instruct as the backbone LLM.
-Considering the ability of LLMs to quickly adapt to downstream tasks with limited data, we followed BIGRec (Bao et al., 2023a) and used relatively smaller datasets compared to traditional models.
-For SFT-based methods, including BIGRec and Re-weighting, the sizes of the training and validation sets were set to $4,096$ and $512$, respectively.
-For DPO-based methods, the sizes were set to $2,048$ and $256$, and training was uniformly conducted on the basis of BIGRec.
-For SPRec, the total number of iterations was set to $5$, with each SFT and DPO phase trained for 1 epoch.
-All experiments were carried out on four RTX 3090 GPUs, each with 24GB of VRAM.
+More details of the implementation are available via https://github.com/RegionCh/SPRec.